OpenStudy (anonymous):

express in terms of i - sqrt -50

OpenStudy (anonymous):

\[\Huge -\sqrt{-50}=-i\sqrt{25\cdot 2}\] can you do it now ? :)

OpenStudy (campbell_st):

use i^2 = -1 \[\sqrt{-50} = \sqrt{2\times 25 \times i^2}\] then \[\sqrt{-50} = 5i \sqrt{2}\]